OpenStudy (anonymous):

how can we calculate the line y=2x+3 on the cartesian plane

OpenStudy (anonymous):

Calculate?

OpenStudy (anonymous):

==

OpenStudy (anonymous):

where is the line y=2x+3 ont eh cartesian plane?

OpenStudy (anonymous):

on the

OpenStudy (anonymous):

u want to plot the line/graph on paper???

OpenStudy (anonymous):

yup because my assessment is tmw and i have to plot it on paper

OpenStudy (anonymous):

to plot it on paper . take a value of x say 1 put it in the equation find the value of y plot the point. take several values of x and find y. join all the points

OpenStudy (anonymous):

does it have to be on the line?

OpenStudy (anonymous):

put values of x in the equation and get corresponding values of y for eg take x=2 y=2x+3 becomes y=2(2)+3 = 4+3 = 7 so u get the pair (2,7) find 3-4 such points, plot them on the graph and join them by a line going a little beyond the extreme points.

OpenStudy (anonymous):

yah thats it.

OpenStudy (anonymous):

k thx

OpenStudy (anonymous):

In case of a linear equation you will get a straight line and all points lying on the line represent the equation

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

y=2(2)+3 y??

OpenStudy (anonymous):

why is this a 2 times 2 here?

OpenStudy (anonymous):

i hv put x=2 in the equation so 2x becomes 2(2) or 2 x 2

OpenStudy (anonymous):

ok...thanks for the answer

OpenStudy (anonymous):

if u take x = 5 then y = 2(5) + 3 = 10 +3 = 13 So if u r satisfied now, pls click GOOD ANSWER button for me :)

OpenStudy (anonymous):

== clicked already...

OpenStudy (anonymous):

many thanks !!

OpenStudy (anonymous):

=o= +.+